OpenStudy (anonymous):

give exact value for: ln(e^2)-log base3.7 1+log base6 9+log base6 4-log10

OpenStudy (anonymous):

\[\Large \ln(e^2)-\log_{3.7}(1)+\log_6(9)+\log_{6}(4)-\log(10)=\] \[\Large 2\ln(e)-0+\log_6(9)+\log_{6}(4)-1=\] \[\Large 2\cdot 1+\log_6(9\cdot 4)-1=\] \[\Large 2+\log_6(36)-1=\] \[\Large 1+\log_6(6^2)=\] \[\Large 1+2\cdot \log_6(6)=\] \[\Large 1+2\cdot 1=3\]

OpenStudy (anonymous):

do you know log rules ?

OpenStudy (anonymous):

i am just learning them. kind of pushed on us at the end of semester and not alot of time to spend on learning them. thank you for the help. i usually learn better if i can see the problem worked out first.

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

TY